# Optimizing Negative Prompts for Enhanced Aesthetics and Fidelity in Text-To-Image Generation

### v1

@@ Abstract
In text-to-image generation, using negative prompts, which describe undesirable image characteristics, can significantly boost image quality. However, producing good negative prompts is manual and tedious. To address this, we propose NegOpt, a novel method for optimizing negative prompt generation toward enhanced image generation, using supervised fine-tuning and reinforcement learning. Our combined approach results in a substantial increase of 25% in Inception Score compared to other approaches and surpasses ground-truth negative prompts from the test set. Furthermore, with NegOpt we can preferentially optimize the metrics most important to us. Finally, we construct Negative Prompts DB, a dataset of negative prompts.
@@ 2 Background
Prompt |
Negative Prompt |
Image |
Other Images |
@@ Appendix B NegOpt Examples
Prompt |
Negative Prompt |
Image |
@@
SFT+RL |
@@
![[Uncaptioned image]](extracted/5465360/assets/girl.jpg)
![[Uncaptioned image]](extracted/5465360/assets/girl_neg_true.jpg)
![[Uncaptioned image]](extracted/5465360/assets/girl_neg_pred.jpg)

### v2

@@ Abstract
In text-to-image generation, using negative prompts, which describe undesirable image characteristics, can significantly boost image quality. However, producing good negative prompts is manual and tedious. To address this, we propose NegOpt, a novel method for optimizing negative prompt generation toward enhanced image generation, using supervised fine-tuning and reinforcement learning. Our combined approach results in a substantial increase of 25% in Inception Score compared to other approaches and surpasses ground-truth negative prompts from the test set. Furthermore, with NegOpt we can preferentially optimize the metrics most important to us. Finally, we detail the construction of Negative Prompts DB, a publicly available dataset of negative prompts.
@@ 2 Background
| Prompt | Negative Prompt | Image | Other Images | |
@@ Appendix B NegOpt Examples
| Prompt | Negative Prompt | Image | |
@@
| SFT+RL | elongated body, cropped image, red eyes, out of frame, draft, deformed hands, signatures, big hair, twisted fingers, double image, long neck, malformed hands, multiple heads, extra limb, ugly… |
@@
![[Uncaptioned image]](extracted/5718673/assets/girl.jpg)
![[Uncaptioned image]](extracted/5718673/assets/girl_neg_true.jpg)
![[Uncaptioned image]](extracted/5718673/assets/girl_neg_pred.jpg)